# Heterogeneous Treatment Effects and

### v3

@@ Abstract
The credibility revolution advances the use of research designs that permit identification and estimation of causal effects. However, understanding which mechanisms produce measured causal effects remains a challenge. A dominant current approach to the quantitative evaluation of mechanisms relies on the detection of heterogeneous treatment effects with respect to pre-treatment covariates. This paper develops a framework to understand when the existence of such heterogeneous treatment effects can support inferences about the activation of a mechanism. We show first that this design cannot provide evidence of mechanism activation without an additional, generally implicit, assumption. Further, even when this assumption is satisfied, if a measured outcome is produced by a non-linear transformation of a directly-affected outcome of theoretical interest, heterogeneous treatment effects are not informative of mechanism activation. We provide novel guidance for interpretation and research design in light of these findings.
Social scientists often make claims about how or why a treatment affects a given outcome. Researchers might tell us, for example, not only that an outreach program increased COVID-19 reporting rates, but that it did so by building trust in government. Or researchers might claim that access to local media increases knowledge of local politics because of the content of news coverage, not because of local political advertising. Claims of this type are often made on the basis of heterogeneous treatment effects (HTEs): the effect of the outreach program on COVID-19 reporting is larger in places that initially distrusted government (Haim, Ravanilla, and Sexton, 2021), for example, or the effect of local media on knowledge of local politics does not change over the election cycle (Moscowitz, 2021). In fact, a majority of recent articles that report HTEs interpret them as evidence about how or why a treatment affects a given outcome.
@@
This framework is important because political scientists so often use HTEs to make claims about causal mechanisms. Surveying the 2021 volumes of three leading journals in political science—the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP)—–we find that a majority (56%) of quantitative studies estimate HTEs and that, conditional on reporting any HTEs, the vast majority of articles (82%) interpret them as providing information about mechanisms. Taken together, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. [^3] Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage. Beyond our analysis, Blackwell, Ma, and Opacic (2024: p. 3) show that HTEs are the modal way that authors in these journals analyze causal mechanisms.
Number of articles: Pr(Report HTEs $\mid$ Pr(Mechanism test $\mid$
Journal (Volume) Total Quant. empirical Reporting HTEs Quant. empirical) Report HTE)
AJPS (65) 61 41 24 0.59 0.83
APSR (115) 102 75 42 0.56 0.90
JoP (83) 142 106 59 0.56 0.76
Total 305 222 125 0.56 0.82
@@
We ask an essential but as-yet-unanswered question: under what conditions do HTEs provide evidence of mechanism activation? We do so by extending the workhorse causal mediation framework (Imai, Keele, and Tingley, 2010). We define a mechanism as an underlying process that influences experience in order to produce a (causal) effect if it is activated (Slough and Tyson, 2024a). In the context of a running example about the effect of adverse shocks like a natural disaster on pro-incumbent voting, voter learning about an incumbent has been posited as a potential mechanism by which such shocks affect voters’ decisions at the ballot box (e.g., Ashworth, Bueno de Mesquita, and Friedenberg, 2018). Whereas the mechanism itself is an unobserved theoretical construct, a mediator is its empirical counterpart: the voter’s posterior belief about the incumbent’s compentence, for example. If the learning mechanism is active, the mediator (a posterior belief) should be affected by the treatment (the negative shock), and it should have a non-zero (indirect) effect on the outcome (vote choice). In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator, or pre-treatment covariate, that is thought to predict the degree to which treatment activates a mechanism and/or the degree to which the mechanism affects on the outcome of interest. In our running example, voters’ prior beliefs about the incumbent’s competence could be a moderator: prior beliefs should condition the extent of learning from the negative shock. Researchers might therefore argue that the existence of HTEs of the adverse shock on pro-incumbent voting at different levels of prior beliefs (the moderator) provide evidence in favor of the voter learning mechanism.
Our results characterize the conditions under which such conclusions are valid, i.e., the conditions under which heterogeneity in conditional average treatment effects (CATEs) with respect to a given moderator is sufficient to show that the mechanism is active. We consider a mechanism to be active when its indirect effect is non-zero for some unit.)
The first such condition is that the moderator of interest is excluded with respect to other (average) effects of the treatment on the outcome. If, for example, voter prior beliefs (the moderator) were to condition the effect of an adverse shock on pro-incumbent voting via mechanisms other than voter learning—such as via directional motives or motivated reasoning (Little, Schnakenberg, and Turner, 2022)—then the exclusion assumption would be violated. The intuition is straightforward: The use of HTEs to assess the activation of a given mechanism relies on the additive separability of the effect of that mechanism from the effects of other mechanisms. In the absence of this exclusion assumption, the relationship between HTEs and the indirect effect of a mechanism is unspecified. This assumption is implicitly invoked by current practice but is rarely explicitly stated or defended.
The second condition concerns the relationship between measured outcomes and the theory. When the exclusion assumption does hold, what can we infer about a mechanism of interest from HTEs with respect to a given moderator (covariate)? Our main results characterize what we can learn from the presence or absence of HTEs. We show that whether such inferences are justified depends on the relationship between measured outcomes and the theory. We thus distinguish between outcome variables that are theorized to be directly affected by a mechanism versus (causally) subsequent outcomes that are indirectly affected.
First, consider the case in which a measured outcome is directly affected by the mechanism of interest. This would be the case if, for example, a researcher was measuring voter utility using an elicitation of voter willingness to pay for the incumbent over the challenger. We show that when the exclusion assumption holds, the existence of HTEs does provide evidence of mechanism activation. This result broadly supports current practices in applied research, albeit while making the underlying assumption explicit. But the absence of HTEs does not rule out mechanism activation. Indeed, the absence of HTEs could arise because: (1) the mechanism is inert (e.g., voters do not learn about the incumbent’s competence from a negative shock); (2) the moderator does not in fact affect the mechanism (e.g., a voter’s prior belief does not condition how much she learns from a shock); and/or (3) because the mechanism produces the same effect on all units (e.g., voter learning produces the same effect on all voters’ assessment of the incumbent). We cannot distinguish between these three possibilities from a lack of HTEs, meaning that we cannot use the absence of HTEs cannot “rule out” a mechanism. This finding invalidates common interpretations of the absence of HTEs in the applied literature.
We further consider the common case in which researchers observe outcomes that are indirectly affected by a mechanism. For example, as is standard in studies of exogenous shocks and voting, researchers measure vote choice. In these cases, a treatment may affect a voter’s (latent) utility, but researchers measure only their choice of incumbent or challenger. When there is a non-linear mapping from the directly affected outcome to the observed, indirectly affected outcome (as in this example), the relationship between HTEs and mechanisms changes. In this case, the existence of HTEs with respect to a moderator no longer provides evidence of mechanism activation, even when the exclusion assumption holds. The logic for this result is straightforward: the non-linear mapping breaks the additive separability of the indirect effect of interest from other indirect and direct effects, which undermines our ability to link HTEs to the indirect effect of a mechanism. This means that even if voters respond to a negative shock only via motivated reasoning—not Bayesian learning—there will generally exist HTE on vote choice when using voters’ prior beliefs as a moderator. This means that the existence of HTEs does not provide evidence of voter learning as a mechanism. Again, this finding counters common interpretations of the presence of HTE for indirectly-affected outcomes.
@@
This paper makes three principal contributions. First, we introduce a new framework to understand the theoretical relationship between causal mechanisms and treatment effect heterogeneity. A large methodological literature provides guidance on the estimation of hetergeneous treatment effects (or “interaction effects”) (e.g., Brambor, Clark, and Golder, 2017; Berry, DeMeritt, and Esarey, 2009; Hainmueller, Mummolo, and Xu, 2018; Grimmer, Messing, and Westwood, 2017; Athey, Tibshirani, and Wager, 2019). More recent contributions the use of HTEs for extrapolation, prediction, or targeting of treatments (e.g., Egami and Hartman, 2022; Huang, 2022; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Yet, because these contributions are primarily statistical, they do not facilitate theoretical analysis of the relationship between mechanisms and the (measured) treatment effects that they produce. Our primary results on the theoretical properties of HTEs—and the problems with existing practice—therefore complement our understanding important statistical issues associated with the estimation of HTEs, namely limited statistical power (e.g., McClelland and Judd, 1993) and multiple-comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink, McConnell, and Vollmer, 2014).
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth, Berry, and Bueno de Mesquita, 2021; Ashworth, Berry, and de Mesquita, 2023; Abramson, Koçak, and Magazinnik, 2022; Slough, 2023). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. This analysis complements recent work by Blackwell, Ma, and Opacic (2024), who make explicit the strong assumptions that underpin the use of treatment effects on intermediate outcomes to discern causal mechanisms. Second, we introduce questions about how measured outcomes relate to theoretical constructs by distinguishing between directly- and indirectly-affected outcomes. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023, 2024a, 2024b), it has not been widely explored in the single-study environment.
@@
Finally, we provide practical guidance for empirical researchers who want to learn about which mechanisms generate observed effects. Our assumption and results reveal a minimal set of attributes of an applied theory that can support the use of HTEs to learn about mechanisms. Two of these attributes, the relationships between (1) a covariate of interest and other mechanisms and (2) measured outcomes and theoretical objects of interest, are generally not discussed in applied work. Second, we show how interpretation of HTEs can be improved, returning to the statistical problems that are well known in this literature. Third, we discuss how our analysis can be used to inform prospective research design. Finally, we show that in order to infer mechanism activation from HTEs in the case of indirectly-observed outcomes, stronger theoretical assumptions about the mapping from directly-observed to indirectly-observed outcomes are necessary. Collectively, these suggestions allow practitioners to accurately use—or, when indicated, avoid—HTEs as a quantitative test of mechanisms.
1 Current Practice
2 Motivating Example: Exogenous Shocks and Voting
Consider a large class of natural experiments on the effect of some exogenous shock, denoted $\omega$, on voter beliefs and behavior in a democracy. [^4] One could alternatively conduct a survey or field experiment in which researchers provide information on an exogenous shock that is otherwise imperfectly observed by voters. This shock could be a natural disaster (e.g., Healy and Malhotra, 2010; Achen and Bartels, 2017), a disease outbreak (e.g., Baccini, Brodeur, and Weymouth, 2021), an economic crisis (e.g., Wolfers, 2002), or even a seemingly-irrelevant event (e.g., Healy, Malhotra, and Mo, 2010). In order to characterize the effect of the shock on voter beliefs and behavior, we adapt a formal model by Ashworth, Bueno de Mesquita, and Friedenberg (2018).
We will assume that an incumbent at the time of the shock is of type $\theta\in\{\underline{\theta},\overline{\theta}\}$, where $\overline{\theta}>\underline{\theta}$, such that a politician of type $\overline{\theta}$ is “good type” and a politician of type $\underline{\theta}$ is a “bad type.” Voters do not observe the politician’s type directly, but may be able to learn about their type from observed governance outcomes. The governance outcome is given by:
$g=f(\theta,\omega)+\varepsilon.$ (1)
In this formulation, higher values of $\omega$ correspond to a more adverse shock (e.g., a more intense natural disaster). Function $f$ is monotonically increasing in $\theta$ and decreasing in $\omega$. $\varepsilon$ is an idiosyncratic shock to the governance outcome that is drawn from a symmetric, continuously differentiable probability density function, $\phi$, that satisfies the monotone likelihood ratio property relative to $g$ [^5] This implies that if $x^{\prime}>x$, then $\frac{\phi(g-x^{\prime})}{\phi(g-x)}$ is strictly increasing in $g$.
Each voter’s utility from a politician depends on the politician’s type, $\theta$, and a valence shock for the incumbent, $v_{i}$. Here, valence captures any attribute of the incumbent that does not depend on their type, including, but not limited to, bias toward a candidate or ideological proximity. Politician type and valence are additively separable, such that voter utility is given by:
$u_{i}=\begin{cases}\theta^{I}+v_{i}&\text{ for the incumbent }(I)\\ \theta^{C}&\text{ for the challenger }(C)\end{cases}$ (2)
In the population, $v_{i}\sim U(-1,1)$. Voters have heterogeneous prior beliefs about the probability that incumbent is of type $\overline{\theta}$, formally $\pi_{i}(\overline{\theta})\in[0,1]$, where $\pi_{i}(\overline{\theta})\sim F_{\pi}$. For simplicity, voters share a common prior $\pi^{C}\in(0,1)$ for the challenger. [^6] The superscript $C$ denotes the challenger. Any belief without a superscript pertains to the incumbent. This simplifying assumption of asymmetry in voters’ priors about the incumbent versus challenger allows us to focus on core related to interpretation of HTEs. However, such a scenario could plausibly be generated by varying attention to the news or governance outcomes during the incumbent’s term (prior to the shock). Voters vote either for the incumbent or challenger. The sequence is as follows:
  1. 1.
    Nature reveals shock $\omega$ and voters observe both the shock and the governance outcome.
  2. 2.
    Voters update their beliefs about the incumbent’s type.
  3. 3.
    Voters vote for either the incumbent or the challenger.
Posterior beliefs and voting behavior are straightforward to characterize. Given a shock, $\omega$ and a voter’s prior, $\pi_{i}$, a voter’s posterior belief about the incumbent’s type, $\theta^{I}$, is given by:
$\beta(\overline{\theta}|\pi_{i},\omega)$ $=\frac{\pi_{i}\phi(g-f(\overline{\theta},\omega))}{\pi_{i}\phi(g-f(\overline{\theta},\omega))+(1-\pi_{i})\phi(g-f(\underline{\theta},\omega))}=\frac{1}{1+\frac{1-\pi_{i}}{\pi_{i}}\frac{\phi(g-f(\underline{\theta},\omega))}{\phi(g-f(\overline{\theta},\omega))}}$ (3)
A voter will vote for the incumbent if and only if:
$\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}\geq\pi^{C}$ (4)

### v4

# C[$\arraybackslashm#1$

@@ Abstract
The credibility revolution advances the use of research designs that permit identification and estimation of causal effects. However, understanding which mechanisms produce measured causal effects remains a challenge. The dominant current approach to the quantitative evaluation of mechanisms relies on the detection of heterogeneous treatment effects (HTEs) with respect to pre-treatment covariates. This paper develops a framework to understand when the existence of such heterogeneous treatment effects can support inferences about the activation of a mechanism. We show first that this design cannot provide evidence of mechanism activation without additional, generally implicit, exclusion assumptions. Further, even when these assumptions are satisfied, the presence of HTEs supports the inference that mechanism is active but the absence of HTEs is generally uninformative about mechanism activation. We provide novel guidance for interpretation and research design in light of these findings.
Social scientists often make claims about how or why a treatment affects a given outcome. Researchers might tell us, for example, not only that an outreach program increased COVID-19 reporting rates, but that it did so by building trust in government. Or researchers might claim that access to local media increases knowledge of local politics because of the content of news coverage, not because of local political advertising. Claims of this type are often made on the basis of heterogeneous treatment effects (HTEs): the effect of the outreach program on COVID-19 reporting is larger in places that initially distrusted government (Haim et al., 2021), for example, or the effect of local media on knowledge of local politics does not change over the election cycle (Moscowitz, 2021). In fact, a majority of recent articles that report HTEs interpret them as evidence about how or why a treatment affects a given outcome.
@@
This framework is important because political scientists so often use HTEs to make claims about causal mechanisms. Surveying the 2021 volumes of three leading journals in political science—the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP)—we find that a majority (56%) of quantitative studies estimate HTEs and that, conditional on reporting any HTEs, the vast majority of articles (82%) interpret them as providing information about mechanisms. Taken together, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. [^3] Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage. Beyond our analysis, Blackwell et al. (2024: p. 3) show that HTEs are the modal way that authors in these journals analyze causal mechanisms.
Number of articles: Prevalence of HTEs Importance
Quantitative Reporting Pr(Report HTEs $\mid$ Pr(Mechanism test $\mid$ Pr(In abstract $\mid$
Journal (Volume) Total empirical HTEs Quant. empirical) Report HTE) Mech. test)
AJPS (65) 61 41 24 0.59 0.83 0.65
APSR (115) 102 75 42 0.56 0.90 0.63
JoP (83) 142 106 59 0.56 0.76 0.82
Total 305 222 125 0.56 0.82 0.72
@@
Clearly, the use of HTEs for mechanism attribution is quite common. However, it may be the case that these tests are viewed as secondary in importance to the “main effect” or the treatment effect in the full sample. While there is likely variation in the extent to which readers value main effects versus HTEs, authors commonly emphasize these mechanism tests as important results. The right column of Table 1 shows that conditional on relying on HTEs for mechanism evaluation, a majority (72%) of articles mention these results about mechanisms in the abstract. Given word constraints, we view this as evidence that authors place non-trivial weight on HTE-based tests of mechanisms as important results of their analyses.
We ask an essential but as-yet-unanswered question: under what conditions do HTEs provide evidence of mechanism activation? We do so by extending the workhorse causal mediation framework (Imai et al., 2010a). We define a mechanism as an underlying process that influences experience in order to produce a (causal) effect when activated (Slough and Tyson, 2024). In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator, or pre-treatment variable, that is thought to predict the degree to which treatment activates a mechanism and/or the degree to which the mechanism affects on the outcome of interest. Our framework provides a minimal structure necessary to link moderators to mechanisms to understand what can be learned from the presence or absence of HTEs. Our results characterize the conditions under which HTEs—specifically, a difference in conditional average treatment effects (CATEs) at different levels of a moderator—are sufficient to show that a specific mechanism is active.
Our framework first elucidates the assumptions that are invoked for mechanism attribution. Specifically, if a covariate moderates the effect of a mechanism, it cannot also be a moderator for any other mechanism(s). If it were, there would be no way to determine which mechanism is responsible for the observed heterogeneity. Our first identification result shows that a difference in CATEs is generically equivalent to the difference in the conditional average indirect effect attributable to a mechanism if and only if exclusion assumptions hold. This means that the covariate does not moderate the effects of other mechanisms. Comparing these assumptions to those invoked by other methods for the quantitative study of mechanisms, namely the assumption of sequential ignorabililty in causal mediation analysis, the exclusion assumptions are neither (logically) stronger nor weaker than sequential ignorability. This means that using HTEs for mechanism detection is not more or less agnostic than mediation analysis. However, one crucial benefit of HTEs is that they can provide information about mechanisms without measurement of mediators (Imai et al., 2011: p. 785).
Learning about a mechanism from HTEs requires more than exclusion assumptions. We introduce the concept of a mechanism detector variable (MDV), a covariate that predicts a stronger activation of a mechanism or a stronger effect of a mechanism on an outcome. Under the exclusion assumptions, the existence of HTEs at different levels of a covariate reveals that the covariate is an MDV. This is, in turn, sufficient to infer that the mechanism in question is active for at least one unit in the sample. This accords with current interpretation of mechanism tests that rely on HTEs. However, if HTEs do not exist for a covariate that is a candidate MDV, we do not learn whether the mechanism is active or inert. The lack of heterogeneity could be a consequence of misspecification of the theory, a mechanism that produces the same effect for all units, or an inert/inactive mechanism. This limits our ability to rule out the activation of a mechanism using HTEs.
Finally, we show how choices about which outcomes to measure and how to measure them can limit the use of HTEs as a test of mechanisms. Specifically, the exclusion assumptions that we articulate imply that the effects of a mechanism on an outcome should be additively separable from the effects of other mechanisms on that outcome. But common non-linear transformations of the outcome variable—e.g., binning, logging, or winsorization—violate this property of additive separability, generating heterogeneity that is not informative about mechanism activation. [^4] Our focus on additive separability stems from the widespread current practice of comparing conditional treatment effects across subgroups. Different measures of mechanistic influence will generally imply different functional form assumptions. This finding suggests that the way that we measure the effects of a mechanism can limit our ability to attribute those very effects to the mechanism. It reveals the need to be more explicit about the relationship between mechanisms and measured outcomes than is common practice.
@@
This paper makes three principal contributions. First, we introduce a new framework to understand the theoretical relationship between causal mechanisms and treatment effect heterogeneity. A large methodological literature provides guidance on the estimation of heterogeneous treatment effects (or “interaction effects”) (e.g., Brambor et al., 2006; Berry et al., 2009; Hainmueller et al., 2018; Grimmer et al., 2017; Athey et al., 2019). More recent contributions the use of HTEs for extrapolation, prediction, or targeting of treatments (e.g., Egami and Hartman, 2022; Huang, 2024; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Yet, because these contributions are primarily statistical, they do not facilitate theoretical analysis of the relationship between mechanisms and the (measured) treatment effects that they produce. Our primary results on the theoretical properties of HTEs—and the problems with existing practice—therefore complement our understanding important statistical issues associated with the estimation of HTEs, namely limited statistical power (e.g., McClelland and Judd, 1993) and multiple-comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink et al., 2014).
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth et al., 2021, 2023; Abramson et al., 2022; Slough, 2023). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. This analysis complements recent work by Blackwell et al. (2024), who make explicit the strong assumptions that underpin the use of treatment effects on intermediate outcomes to discern causal mechanisms. Second, we introduce questions about how measured outcomes relate to theoretical constructs. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023, 2024, 2025), it has not been widely explored in the single-study environment.
@@
Finally, we provide practical guidance for empirical researchers who want to learn about which mechanisms generate observed effects. We illustrate this guidance concretely by analyzing four recent empirical studies about how partisan affinity (or bias) condition voter responses to corruption revelation (Anduiza et al., 2013; Arias et al., 2019; de Figueiredo et al., 2023; Eggers, 2014). Our assumptions and results reveal a minimal set of attributes of an applied theory that can support the use of HTEs to learn about mechanisms. Two of these attributes, the relationships between (1) a covariate of interest and other mechanisms and (2) measured outcomes and theoretical objects of interest, are generally not discussed in applied work. Second, we show how interpretation of HTEs can be improved, returning to the statistical problems that are well known in this literature. Third, we discuss how our analysis can be used to inform prospective research design. Finally, we consider the merits of invoking additional assumptions or statistical models to address some of the issues we identify. Collectively, these suggestions allow practitioners to accurately use—or, when indicated, avoid—HTEs as a quantitative test of mechanisms.
1 Motivating Example: Corruption Revelation and Voting
Despite widespread unpopularity of corruption by public officials in public opinion polls, voters in many democracies routinely re-elect politicians who engage in corruption. One explanation for the prevalence of public corruption in many democracies is that corruption by specific politicians is not observed by voters. If voters were to receive information that an incumbent or candidate were corrupt, then they should exhibit less support for the candidate. Scholars have evaulated the empirical merits of this argument using field experiments (e.g., Dunning et al., 2019), randomized corruption audits (e.g., Ferraz and Finan, 2008), survey experiments (e.g., Anduiza et al., 2013), and observational studies (e.g., Eggers, 2014). Existing meta-studies document heterogeneity in voter responses to such information (Incerti, 2020; Slough, 2024). Here, we examine whether such heterogeneity is informative about a voter (dis)taste for corruption mechanism.
Specifically, consider a randomized experiment in which the treatment reveals that an incumbent in a constituency has engaged in corruption. We will say that $c_{i}=1$ if voter $i$ receives the corruption information treatment and $c_{i}=0$ if they do not receive the information. Voters value multiple attributes of politicians. First, they dislike corruption, albeit to different degrees, where $\lambda_{i}\in(0,1)$ captures variation in corruption aversion across voters. Second, voters value partisan alignment with politicians. Specifically, voter $i$ may be aligned $(a_{i}=1)$, independent/neutral ($a_{i}=0$), or unaligned $(a_{i}=-1)$ with an incumbent politician. Finally, voters idiosyncratically value valence characteristics of a politician (e.g., their personality or personal background), which we represent with the random variable $\varepsilon_{i}$. We assume that $\varepsilon_{i}$ is independent of $\lambda_{i}$ and $a_{i}$, and follows a standard normal distribution. Each voter’s utility from a vote for the incumbent is given by:
$u_{i}$ $=-\lambda_{i}c_{i}+a_{i}+\varepsilon_{i}$
Corruption information enters voters’ assessment of the incumbent through their distaste for corruption, $\lambda_{i}$. One can think of the product $\lambda_{i}c_{i}$ as akin to a mediator that captures the degree to which this distaste is activated by the information treatment $c_{i}$. Ultimately voters choose between the incumbent and a challenger. Without loss of generality, we normalize a voter’s utility from a vote for the challenger to be 0. Therefore, voters will vote for the incumbent if $u_{i}\geq 0$.